# Affine Frequency Division Multiplexing (AFDM) for 6G: Properties, Features, and Challenges

### v1

@@ I Introduction
Enabling such advanced applications and satisfying such stringent requirements demand, however, wireless systems to operate in increasingly higher frequency bands, such as the millimeter wave (mmWave) and Terahertz (THz) spectra, subject to multipath and high-mobility. This gives rise to doubly-dispersive channels that exhibit both time- and frequency-selectivity [5], leading to severe deterioration of received signals and, consequently, substantial challenges for communications systems employing conventional waveforms [6]. In particular, the orthogonal frequency division multiplexing (OFDM) waveform, widely used in fourth generation (4G) and fifth generation (5G) systems, struggles in such environments due to severe inter-symbol interference (ISI) and inter-carrier interference (ICI) caused by multipath delays and especially Doppler shifts, which disrupts OFDM subcarrier orthogonality.
Furthermore, 6G is anticipated to bring a paradigm shift towards a native, and hence, optimized, support of ISAC [2], enabling a wide range of multifunctional services, real-time situational awareness, improved spectrum utilization, and high-resolution localization [7]. These features are not easily available through conventional waveforms, which are susceptible to Doppler interference and prevent accurate velocity estimation,estimation, thus motivating the design of novel waveforms that can offer strong Doppler robustness and high-resolution sensing capabilities.
@@
Another recent proposal is affine frequency division multiplexing (AFDM), which has emerged as a promising candidate for 6G [10], whereby data based is modulated via the affine Fourier transform (AFT), an affine generalization of the conventional linear Fourier transform (FT) employed in OFDM. By leveraging two tunable parameters that control the twistedness of the affine time-frequency domain, AFDM is capable of adapting to the specific characteristics of doubly-dispersive channels, enabling optimal diversity gain and delay-Doppler orthogonality. In addition, these tunable parameters facilitate a wide range of waveform design opportunities, supporting inherent advanced functionalities, such as index modulation (IM), peak-to-average power ratio (PAPR) reduction, and physical-layer security [11].
@@
Within ongoing discussions on candidate waveforms for 6G, a key point of debate among industry, academia, and standardization bodies such as the 3rd Generation Partnership Project (3GPP), is whether 6G necessitates entirely new waveform paradigms or refinements of existing ones. Rather than replacing legacy waveforms like OFDM, many advocate for expanding the waveform toolbox to enable flexible, context-aware operation, highlighting the need for candidates that balance performance, compatibility, and deployability. In this context, AFDM stands out for its inherent compatibility with OFDM-based systems, enabled by its generalization of the FT to the AFT, facilitating seamless integration and coexistence with existing 4G and 5G infrastructure [12].
@@
All in all, the above collectively set AFDM as a strong contender for standardization in the 6G era. Therefore, we aim at providing a comprehensive tutorial on, and motivation for AFDM, exploring its fundamental properties and distinctive features, as well as challenges and opportunities for its adoption in 6G networks. The article is structured as follows: Section II details the fundamentals and key properties of the AFDM waveform; Section III examines eight of the advantageous features of AFDM and evaluates their implications to relevant 6G applications; Section IV then addresses another set of eight points on opportunities, challenges, and pathways towards standardization of AFDM in future 6G systems; and finally, Section V concludes with a summary and prospects.
@@ II Fundamentals and Key Properties
Figure 1: Illustrative comparison of OFDM, OCDM, AFDM, and OTFS waveforms in terms of: a) the modulator structure in terms of the core Fourier transform and the relationship of the chirp frequencies $c_{1}$ and $c_{2}$, which shows that AFDM generalizes OFDM and OCDM, b) the time-domain subcarriers (real component only), c) the time-frequency occupancy of the subcarriers, d) effective channel structure with four arbitrary integer delay-Doppler taps, and e) the same channels of d) with fractional Doppler shifts, causing Doppler-domain leakage.
In this section, we briefly introduce the fundamentals of the AFDM signal modulation, highlighting the inherent properties of the waveform, providing a background towards the consequent features and applications to be elaborated in the following sections. Given the breadth of the audience and the tutorial nature of the IEEE Magazine, this article presents an illustrative and intuitive overview of the AFDM fundamentals, omitting detailed derivations. Readers interested in the full theoretical formulation and rigorous analysis of the AFDM are encouraged to consult [10, 6].
@@ II-A Signal Generation and Detection
Furthermore, the frequency sweep rate of the chirp subcarriers in AFDM is fully configurable via the two tunable chirp parameters inherent to the IDAFT (often denoted by $c_{1}$ and $c_{2}$). As will be further elaborated upon later, this unique design flexibility through the inclusion of tunable chirp parameters - not present in contending waveforms - is a key feature of AFDM, which ensures orthogonality and full diversity in doubly-dispersive channels, in addition to enabling a plethora of waveform optimization and applications.
@@
Finally, before transmission, the AFDM signal must be prepended with a chirp-periodic prefix (CPP), which while structurally similar to the conventional cyclic prefix (CP) in OFDM, includes deterministic phase adjustments determined by the chirp parameters. With the CPP, AFDM satisfies the quasi-periodicity conditions and ensures the circular convolution compatibility to preserve the symbol orthogonality.
At the receiver, demodulation is performed using a matched forward discrete affine Fourier transform (DAFT), which is configured with the same chirp parameters $c_{1}$ and $c_{2}$ used in the transmitter-side IDAFT, ensuring coherent symbol recovery. For appropriately selected chirp parameters, whose optimal values can be obtained in closed-form based on the channel statistics (maximum unambiguous delay and Doppler shift), and a desired guard width, the effective channel matrix becomes a structured superposition of shifted band-diagonal components. Each such component corresponds to a distinct delay-Doppler path, and their orthogonality minimises ISI and ICI even under significant channel dispersion, further ensured by the selected guard width parameter.
II-B Affine Fourier Domain and the Evolution of 6G Waveforms
Chirp signals are well-known to have favorable properties under both time shifts (delays) and frequency shifts (Doppler), which intuitively suggests them to be well-matched to channels where such effects are jointly present, i.e., the doubly-dispersive channels. As described in the previous section, AFDM exploits this notion, and modulates the data symbols unto the orthogonal chirp subcarriers via the inverse discrete affine Fourier transform (IDAFT).
Trivially, the IDAFT originates from the affine Fourier transform (AFT), also referred to as the linear canonical transform (LCT) [13], a family of unitary linear operators defined by four configurable parameters $(a,b,c,d)$, satisfying $ad-bc=1$. By selecting appropriate values of the parameters $a,b,c,$ and $d$, the AFT generalizes a wide range of well-known and fundamental transforms as its special cases.
For example, the AFT is reduced to the conventional Fourier transform (FT)with fixed parameters $(a,b,c,d)=(0,1,-1,0)$, while the fractional Fourier transform (FrFT)generalizes this by introducing a tunable angle parameter $\alpha\in[0,2\pi)$ such that $(a,b,c,d)=(\cos\alpha,\sin\alpha,-\sin\alpha,\cos\alpha)$. On the other hand, the Fresnel transform (FnT), associated with free-space wave propagation and used in the modulation of OCDM, is obtained when $(a,b,c,d)=(1,z,0,1)$, where $z$ is a tunable parameter representing the propagation distance.
@@ II-B Affine Fourier Domain and the Evolution of 6G Waveforms
[…]
@@ III-A Delay-Doppler Robustness and Diversity
The AFDM waveform stands out in this regard over other waveforms in the communications performance, as illlustrated in Fig. 2, offering inherent robustness to doubly-dispersive channels with arbitrary delay-Doppler statistics. Through the parametrizable chirp subcarriers, AFDM has been shown to guarantee full diversity without the need for precoding or equalization enhancements, by optimizing the first chirp parameter $c_{1}$, for which a closed-form solution is available [10].
This contrasts with conventional OFDM, where each symbol is confined to a single time-frequency subcarrier, resulting in a maximum diversity order of one in doubly-dispersive channels. While OTFS improves over the OFDM by spreading each symbol over the entire time-frequency grid via its delay-Doppler representation and transforms, uncoded OTFS has been shown to fall short of full diversity in the high-signal-to-noise ratio (SNR) regime unless mitigation strategies such as delay-Doppler precoding are applied [14].
@@
Figure 2: BER performance of OFDM, OCDM, OTFS, and AFDM over a doubly-dispersive channel with $P=3$ paths. System parameters are: number of subcarriers $N=64$, constellation cardinality $M=4$, OTFS grid size $K\times L=8\times 8$, carrier frequency $f_{c}=50$GHz, bandwidth $B=150$MHz, AFDM guard parameter $\xi=1$, AFDM pre-chirp parameter $c_{2}=1/{2\pi N}$.
@@ III-C CSI-based Radar Parameter Estimation
Fundamentally, the full delay-Doppler representation of the doubly-dispersive channel is very closely linked to the radar parameters of environmental scatterers. Namely, each multipath component is associated with its delay and Doppler shift, which correspond to the bistatic range and velocity of the scatterer (or target). This duality has motivated a wide range of methods for extracting such physical parameters from the estimated channel state information (CSI), or conversely to perform channel estimation directly via parameterized models that exploit the deterministic structure of the delay-Doppler channel, which has shown great success and efficiency [15].
@@ III-D Conventional Radar Processing and Ambiguity Function
While CSI-based radar parameter extraction has become central to many modern communication-centric sensing frameworks, it is equally critical to consider the classical signal processing foundation of radar sensing: correlation-based matched filtering.
This classical perspective is especially relevant for AFDM, whose underlying chirp subcarriers strongly resemble the gold standard of chirp structures historically used in radar systems. In this context, the ambiguity function (AF) is fundamental to quantify how effectively a probing waveform can resolve targets in the delay and Doppler dimensions, characterizing the resolution and sidelobe suppression.
As illustrated in Fig. 3, in contrast to the poor ambiguity properties of OFDM in the Doppler domain, suffering from severe Doppler smearing and limited velocity resolution, or similarly of the OCDM in the delay-domain, AFDM exhibits a sharply localized, spike-like ambiguity function in both of the delay and Doppler dimensions, similarly to the OTFS, but has additionally shown to provide delay-Doppler lattice alignment options via the configurable of the chirp parameters.
(a) Zero-Doppler cut.
(b) Zero-delay cut.
Figure 3: Normalized ambiguity functions of OFDM, OCDM, OTFS, and AFDM in the delay-Doppler domain, with unitary sensing symbols. System parameters are: number of subcarriers $N=64$, OTFS grid size $K\times L=8\times 8$, carrier frequency $f_{c}=50$GHz, bandwidth $B=150$MHz, sampling frequency $f_{s}=2B$ Hz, AFDM guard parameter $\xi=1$, AFDM pre-chirp parameter $c_{2}=1/{2\pi N}$.
@@ III-E AFDM for Integrated Sensing and Communciations (ISAC)
Indeed, recent works have shown that AFDM achieves accurate delay and Doppler parameter recovery with minimal sensing overhead but without sacrificing communication performance, leveraging embedded pilot modulation, subcarrier-level index modulation, and bilinear estimation methods for joint channel and data recovery [15].
@@ III-F Frequency-Domain Processing
As expected from the close relationship between the AFDM and OFDM, for specific values of the chirp parameters $c_{1}$ and $c_{2}$, and certain subcarrier lengths $N$ still achieving full diversity and subcarrier orthogonality - closed-form frequency-domain expressions of the AFDM waveform are available, enabling straightforward implementation and conventional analysis [12]. In other words, AFDM can be viewed as a precoded OFDM system, where the precoder maps the data from the DAFT domain into the frequency domain.
This compatibility with the frequency-domain processing of OFDM is a key advantage of AFDM, as it allows for the direct application of established OFDM techniques, such as frequency-domain pilot placement, channel estimation and equalization, and digital beamforming.
Furthermore, this also carries an important implication that a smooth transition from 5G to 6G, i.e., from OFDM to AFDM, can be achieved without major modifications to existing hardware and infrastructure. This makes AFDM a practical and backward-compatible candidate for future wireless systems, as further discussed in Section IV-A.

### v3

@@ I Introduction
Enabling such advanced applications and satisfying such stringent requirements demand, however, wireless systems to operate in increasingly higher frequency bands, such as the millimeter wave (mmWave) and Terahertz (THz) spectra, subject to multipath and high-mobility. This gives rise to doubly-dispersive channels that exhibit both time- and frequency-selectivity, leading to severe deterioration of received signals and, consequently, substantial challenges for communications systems employing conventional waveforms [5]. In particular, the orthogonal frequency division multiplexing (OFDM) waveform, widely used in fourth generation (4G) and fifth generation (5G) systems, struggles in such environments due to severe inter-symbol interference (ISI) and inter-carrier interference (ICI) caused by multipath delays and especially Doppler shifts, which disrupts OFDM subcarrier orthogonality.
Furthermore, 6G is anticipated to bring a paradigm shift towards a native, and hence, optimized, support of ISAC [2], enabling a wide range of multifunctional services, real-time situational awareness, improved spectrum utilization, and high-resolution localization [6]. These features are not easily available through conventional waveforms, which are susceptible to Doppler interference and prevent accurate velocity estimation, thus motivating the design of novel waveforms that can offer strong Doppler robustness and sensing capabilities.
@@
Another recent proposal is affine frequency division multiplexing (AFDM), which has emerged as a promising candidate for 6G [9], whereby data based is modulated via the affine Fourier transform (AFT), an affine generalization of the conventional linear Fourier transform (FT) employed in OFDM. By leveraging two tunable parameters that control the twistedness of the affine time-frequency domain, AFDM is capable of adapting to the specific characteristics of doubly-dispersive channels, ensuring optimal diversity gain – i.e., exploit multiple independently fading paths with varying time-frequency selectivity for improved reliability. In addition, these tunable parameters facilitate a wide range of waveform design opportunities, supporting inherent advanced functionalities, such as index modulation (IM), peak-to-average power ratio (PAPR) reduction, and physical-layer security.
Owing to its inherent delay-Doppler localization, directly related to the range and velocity sensing capabilties of the waveform, recent studies have also identified AFDM as a highly effective waveform for ISAC which can well address the communications-sensing trade-off[10]. In such, AFDM offers two fundamental ISAC approaches: a radar-like correlation approach, where the favorable ambiguity function (AF) of AFDM in both delay and Doppler domains enables high range-velocity resolution [11], and another approach based on the deterministic structure of the AFDM effective channel, whose delay-Doppler representation allows channel estimation to be equivalent to the radar parameter estimation, or vice versa [12]. Both approaches have demonstrated strong performances and highlight the potential of AFDM-based ISAC in 6G.
@@
Amidst the research, there are also ongoing discussions on candidate waveforms for 6G, where a key point of debate among industry, academia, and standardization bodies, such as the 3rd Generation Partnership Project (3GPP) and European Telecommunications Standards Institute (ETSI), continues to be whether 6G necessitates entirely new waveform paradigms or refinements of existing ones. In other words, rather than replacing legacy waveforms like OFDM, many advocate for expanding the waveform toolbox to enable flexible, context-aware operation, highlighting the need for candidates that balance performance, compatibility, and deployability.
Indeed, as of October 2025, several companies within the 3GPP are actively proposing new waveform candidates, among which AFDM has been highlighted as a promising option [13]. This interest stems from its inherent compatibility with legacy waveforms such as CP-OFDM and DFT-s-OFDM, which enables seamless integration and coexistence with existing 4G/5G infrastructure, as well as from its channel adaptability – an essential feature for the diverse use cases and deployment scenarios envisioned for 6G [14].
All in all, the above establishes AFDM as a strong candidate waveform for 6G. Accordingly, this article offers a comprehensive and standardization-oriented study of AFDM, examining its key properties, advantages, and challenges specifically in relation to the 6G requirements and deployment considerations. This perspective well complements other studies, for example which focus on AFDM’s waveform characteristics and flexibility across various channel and application scenarios [14], in all providing a holistic view of the waveform’s potential and practical relevance.
@@
The article is structured as follows: Section II details the fundamentals and key properties of the AFDM waveform; Section III examines eight of the advantageous features of AFDM and evaluates their implications to relevant 6G applications; Section IV then addresses another set of eight points on opportunities, challenges, and pathways towards standardization of AFDM in future 6G systems; and finally, Section V concludes with a summary and prospects.
@@ II Fundamentals and Key Properties
In this section, we briefly introduce the fundamentals of the AFDM signal modulation, highlighting the inherent properties of the waveform, providing a background towards the consequent discussions to be elaborated in the following sections [^1] Given the breadth of the audience and the nature of the IEEE Magazine, we present an illustrative and intuitive overview of the AFDM fundamentals, omitting detailed derivations. Readers interested in the full theoretical formulation and analysis of AFDM are encouraged to consult [9, 5]..
@@ II-A Signal Generation and Detection
Figure 1: Illustrative comparison of OFDM, OCDM, AFDM, and OTFS waveforms in terms of: (a) the modulator structure in terms of the core Fourier transform and the relationship of the chirp frequencies $c_{1}$ and $c_{2}$, which shows that AFDM generalizes OFDM and OCDM, (b) the time-domain subcarriers, (c) the time-frequency (delay-Doppler) occupancy of the subcarriers, (d) effective channel structure with four arbitrary integer delay-Doppler taps, and (e) the same channels of (d) with fractional Doppler shifts (i.e., normalized Doppler frequencies do not align perfectly with the discrete integer grid used in delay-Doppler signal processing), causing Doppler-domain energy leakage.
@@
Furthermore, the frequency sweep rate of the chirp subcarriers in AFDM is fully configurable via the two tunable chirp parameters inherent to the IDAFT (often denoted by $c_{1}$ and $c_{2}$), namely by $c_{1}$. As will be further elaborated upon later, this unique design flexibility through the inclusion of tunable chirp parameters - not present in contending waveforms - is a key feature of AFDM, which ensures orthogonality and full diversity in doubly-dispersive channels, in addition to enabling a plethora of waveform optimization and applications.
@@
Finally, before transmission, the AFDM signal must be prepended with a chirp-periodic prefix (CPP), which while structurally similar to the conventional cyclic prefix (CP) in OFDM, includes deterministic phase adjustments determined by the chirp parameters. With the CPP, AFDM satisfies the quasi-periodicity conditions and ensures the circular convolution compatibility to preserve the symbol orthogonality. At the receiver, demodulation is performed using a matched forward discrete affine Fourier transform (DAFT), which is configured with the same chirp parameters $c_{1}$ and $c_{2}$ used in the transmitter-side IDAFT, ensuring coherent symbol recovery.
For appropriately selected chirp parameters, whose optimal values can be obtained in closed-form based on the channel statistics (maximum unambiguous delay and Doppler shift), and a desired guard width, the effective channel matrix becomes a structured superposition of shifted band-diagonal components. Each such component corresponds to a distinct delay-Doppler path, and their orthogonality minimises ISI and ICI even under significant channel dispersion, further ensured by the selected guard width parameter.
B. Affine Fourier Domain and the Evolution of 6G Waveforms
Chirp signals are well-known to have favorable properties under both time shifts (delays) and frequency shifts (Doppler), xwhich intuitively suggests them to be well-matched to channels where such effects are jointly present, i.e., the doubly-dispersive channels. As described in the previous section, AFDM exploits this notion, and modulates the data symbols unto the orthogonal chirp subcarriers via the inverse discrete affine Fourier transform (IDAFT).
Trivially, the IDAFT originates from the affine Fourier transform (AFT), also referred to as the linear canonical transform (LCT), a family of unitary linear operators defined by four configurable parameters $(a,b,c,d)$, satisfying adbc=1ad-bc=1. By selecting appropriate values of the parameters $a,b,c,$ and $d$, the AFT generalizes a wide range of well-known and fundamental transforms as its special cases.
For example, the AFT is reduced to the conventional Fourier transform (FT) with fixed parameters (a,b,c,d)=(0,1,1,0)(a,b,c,d)=(0,1,-1,0), while the fractional Fourier transform (FrFT) generalizes this by introducing a tunable angle parameter $\alpha\in[0,2\pi)$ such that $(a,b,c,d)=(\cos\alpha,\sin\alpha,-\sin\alpha,\cos\alpha)$. On the other hand, the Fresnel transform (FnT), associated with free-space wave propagation and used in the modulation of OCDM, is obtained when $(a,b,c,d)=(1,z,0,1)$, where $z$ is a tunable parameter representing the propagation distance.
@@ B. Affine Fourier Domain and the Evolution of 6G Waveforms
All in all, the AFT may be regarded as a generalized class of linear transforms that unifies and extends classical time-frequency operators, and AFDM as a natural generalization in the evolution of multicarrier modulation schemes, subsuming earlier schemes like OFDM and OCDM as special cases by introducing tunable chirp parameters $c_{1}$ and $c_{2}$, corresponding to the four AFT parameters $(a,b,c,d)$ [9].
@@ III-A Delay-Doppler Robustness and Diversity
The AFDM waveform stands out in this regard over other waveforms in the communications performance, as illlustrated in Fig. 2, offering inherent robustness to doubly-dispersive channels with arbitrary delay-Doppler statistics. The waveforms use the same total number of symbol resources, with OFDM, OCDM, and AFDM waveforms each employing $N$ time-frequency subcarriers, while OTFS operates over a delay-Doppler grid of size $K\times L$, where $K$ and $L$ denote the numbers of delay and Doppler bins, respectively, and $KL=N$.
Through the parametrizable chirp subcarriers, AFDM has been shown to guarantee full diversity without the need for precoding or equalization enhancements, by optimizing the first chirp parameter $c_{1}$, for which a closed-form solution is available [9]. This contrasts with conventional OFDM, where each symbol is confined to a single time-frequency subcarrier, resulting in a maximum diversity order of one in doubly-dispersive channels. While OTFS improves over the OFDM by spreading each symbol over the entire time-frequency grid via its delay-Doppler representation and transforms, uncoded OTFS has been shown to fall short of full diversity in the high-signal-to-noise ratio (SNR) regime unless mitigation strategies such as delay-Doppler precoding are applied [15].
@@
Figure 2: BER performance of OFDM, OCDM, OTFS, and AFDM over a doubly-dispersive channel with $P=3$ paths. System parameters are: number of subcarriers $N=64$, constellation cardinality $M=4$, OTFS grid size $K\times L=8\times 8$, carrier frequency $f_{c}=50$GHz, bandwidth B=150B=150MHz, AFDM guard parameter $\xi=1$, AFDM pre-chirp parameter $c_{2}=\frac{1}{2N}$.
@@ III-C CSI-based Radar Parameter Estimation
Fundamentally, the full delay-Doppler representation of the doubly-dispersive channel is very closely linked to the radar parameters of environmental scatterers. Namely, each multipath component is associated with its delay and Doppler shift, which correspond to the bistatic range and velocity of the scatterer (or target).
This duality has motivated a wide range of methods for extracting such physical parameters from the estimated channel state information (CSI), or conversely to perform channel estimation directly via parameterized models that exploit the deterministic structure of the delay-Doppler channel, which has shown great success and efficiency [10, 12].
@@ III-D Conventional Radar Processing and Ambiguity Function
While CSI-based radar parameter extraction has become central to many modern communication-centric sensing frameworks, it is equally critical to consider the classical signal processing foundation of radar sensing: correlation-based matched filtering. This classical perspective is especially relevant for AFDM, whose underlying chirp subcarriers strongly resemble the gold standard of chirp structures historically used in radar systems. In this context, the AF is fundamental to quantify how effectively a probing waveform can resolve targets in the delay and Doppler dimensions, characterizing the resolution and sidelobe behavior of the waveforms.
Figure 3 illustrates the four waveforms using a fractionally interpolated (for a high-resolution representation of the continuous behavior) AF computed under unitary symbol modulation – i.e., all symbols are set to one. This highlights the intrinsic waveform- and transform-dependent characteristics in the delay-Doppler domain, independent of the behavior introduced by modulated random data symbols [11].
The OFDM waveform exhibits a sharp peak along the delay axis but poor ambiguity characteristics in the Doppler domain, resulting in significant Doppler smearing and limited velocity resolution. In contrast, OTFS demonstrates improved joint localization in both delay and Doppler dimensions, albeit with broader mainlobes. Meanwhile, OCDM and AFDM achieve sharply concentrated, spike-like ambiguity functions in both delay and Doppler, reflecting their inherent chirp-based structure. In addition, AFDM is known to offer flexibility through the configuration of its chirp parameters, enabling delay-Doppler lattice alignment and controllable AF shaping.
(a) Zero-Doppler cut (delay response) of the waveform AFs.
(b) Zero-delay cut (Doppler response) of the waveform AFs.
Figure 3: Interpolated (high-resolution) normalized ambiguity functions of OFDM, OCDM, OTFS, and AFDM in the delay-Doppler domain, with unitary sensing symbols, i.e., all symbols are set to one. System parameters are: number of subcarriers $N=64$, OTFS delay-Doppler grid size $K\times L=8\times 8$, AFDM chirp parameters $c_{1}=\frac{11}{2N},c_{2}=\frac{1}{2N}$.
@@ III-E AFDM for Integrated Sensing and Communciations (ISAC)
Indeed, recent works have shown that AFDM achieves accurate delay and Doppler parameter recovery with minimal sensing overhead but without sacrificing communication performance, leveraging embedded pilot modulation, subcarrier-level index modulation, and bilinear estimation methods for joint channel and data recovery [12, 10].
@@ III-F Frequency-Domain Processing
Given the close structural relationship between AFDM and OFDM, certain configurations of the chirp parameters $(c_{1},c_{2})$ and subcarrier length $N$, while still preserving full diversity and subcarrier orthogonality, admit closed-form frequency-domain representations of the AFDM waveform, enabling straightforward implementation and conventional analysis [16]. From this perspective, AFDM can be interpreted as a precoded OFDM system, where the precoder maps data symbols from the DAFT domain into the frequency domain through chirp-dependent rotations.
The trivial case for $c_{1}=0$, AFDM degenerates to standard OFDM up to a symbol-wise phase rotation and all OFDM-based frequency-domain techniques can be leveraged directly. However for $c_{1}\neq 0$, the modulation introduces additional phase rotations before OFDM-like transformation. Still, the fundamental one-dimensional processing chain, where a complex symbol vector undergoes an DFT followed by another phase rotation, remains structurally identical to that of OFDM.
Accordingly, in non-doubly-dispersive channels (i.e., those exhibiting multipath delay but negligible Doppler), existing frequency-domain techniques developed for OFDM– such as pilot placement, channel estimation, and equalization – can be directly applied to AFDM. This strong structural alignment between the AFDM and OFDM processing chains implies that existing frequency-domain modules can be adapted with minimal modification for specific scenarios [14], facilitating a seamless evolution from 5G to 6G.